Let T be a continuous random variable with pdf

f(x) = (1/10(e)-t/10 , : t ⩾ 0

(0, : x < 0


Define the continuous random variable by

(100, : 0 < T ⩽ 1

X= (50, : 1 < T ⩽ 3

(0, : T > 3

Find E(X)


1
Expert's answer
2021-09-24T06:41:02-0400

we can use "\u222b^\u221e_{\n\u2212\u221e}\nfX(u)du=1:"

"1 \t=\u222b^\u221e_{\u2212\u221e}fX(u)du\\\\\n=\u222b^1_{\u22121}cu^2du\\\\\n=\\frac{2}{3}c."

"EX \t=\u222b^1_{\u22121}uf_X(u)du\\\\\n=\\frac{3}{2}\u222b^1_{\u22121}u^3du\\\\\n=0."